OpenStudy (anonymous):

Suppose 2 moles of an ideal gas are enclosed in a container of volume 2.10·10-4 m3. The container is then placed in a furnace and raised to a temperature of 392 K. What is the final pressure of the gas?

OpenStudy (anonymous):

first use the equivalencies of P.V=n.R.T thus P = n.R.T/V

OpenStudy (anonymous):

P = 2.0,082.392/ 2.10^-4 P = 64,288 . 10^4/2 P = 32,144 . 10^4 atm